# A Library of LLM Intrinsics for Retrieval-Augmented Generation

### v1

@@ Appendix A Technical Appendices and Supplementary Material
A.1 Query Rewrite Quickstart Example
The following code describes how to use the Query Rewrite model.
1
2import torch
3from transformers import AutoTokenizer, AutoModelForCausalLM
4from peft import PeftModel
5import json, re
6
7INSTRUCTION_TEXT = "Reword the final utterance from the USER into a single utterance that doesn’t need the prior conversation history to understand the user’s intent. If the final utterance is a clear and standalone question, please DO NOT attempt to rewrite it, rather output the last user utterance as is. "
8JSON = "Your output format should be in JSON: { \"rewritten_question\": <REWRITE> }"
9REWRITE_PROMPT = "<|start_of_role|>rewrite: " + INSTRUCTION_TEXT + JSON + "<|end_of_role|>"
10
11device=torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’)
12
13BASE_NAME = "ibm-granite/granite-3.2-8b-instruct"
14LORA_NAME = "ibm-granite/granite-3.2-8b-lora-rag-query-rewrite"
15
16tokenizer = AutoTokenizer.from_pretrained(BASE_NAME, padding_side=’left’,trust_remote_code=True)
17model_base = AutoModelForCausalLM.from_pretrained(BASE_NAME, device_map =’auto’)
18model_rewrite = PeftModel.from_pretrained(model_base, LORA_NAME)
19
20# Input conversation
21conv = [
22 {
23 "role":"user",
24 "content":"Tim Cook is the CEO of Apple Inc."
25 },
26 {
27 "role":"assistant",
28 "content":"Yes, Tim Cook is the Chief Executive Officer of Apple Inc."
29 },
30 {
31 "role":"user",
32 "content":"and for Microsoft?"
33 }
34]
35
36# Generate the query rewrite for the last turn in the above conversation
37conv = [{"role":"system", "content":""}] + conv
38input_text = tokenizer.apply_chat_template(conv, tokenize=False) + REWRITE_PROMPT
39inputs = tokenizer(input_text, return_tensors="pt")
40output = model_rewrite.generate(inputs["input_ids"].to(device), attention_mask=inputs["attention_mask"].to(device), max_new_tokens=80)
41output_text = tokenizer.decode(output[0])
42
43# Regex pattern to extract the JSON with the rewrite from the output of the model
[…]
45match_js = re.findall(pattern, output_text)[0]
46try:
47 #Parse the JSON and extract the rewrite
48 rewrite = json.loads (match_js) [’rewritten_question’]
49except Exception as e:
50 rewrite = match_js.split ("\"rewritten_question\": ", 1)[1]
51
[…]
53# Rewrite: Who is the CEO of Microsoft?
A.2 Uncertainty Quantification Quickstart Example
The following code describes how to use the Uncrtainty Quantification model to answer questions and obtain intrinsic calibrated certainty scores. Note that a generic system prompt is included, this is not necessary and can be modified as needed.
1
2import torch,os
3from transformers import AutoTokenizer, AutoModelForCausalLM
4from peft import PeftModel, PeftConfig
5
6token = os.getenv("HF_MISTRAL_TOKEN")
7BASE_NAME = "ibm-granite/granite-3.2-8b-instruct"
8LORA_NAME = "ibm-granite/granite-3.2-8b-lora-uncertainty"
9device=torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’)
10
11# Load model
12tokenizer = AutoTokenizer.from_pretrained(BASE_NAME,padding_side=’left’,trust_remote_code=True, token=token)
13model_base = AutoModelForCausalLM.from_pretrained(BASE_NAME,device_map="auto")
14model_UQ = PeftModel.from_pretrained(model_base, LORA_NAME)
15
16question = "What is IBM Research?"
17print("Question:" + question)
18question_chat = [
19 {
20 "role": "user",
21 "content": question
22 },
23]
24
25# Generate answer with base model
26input_text = tokenizer.apply_chat_template(question_chat,tokenize=False,add_generation_prompt=True)
27
28
29#tokenize
30inputs = tokenizer(input_text, return_tensors="pt")
31output = model_base.generate(inputs["input_ids"].to(device), attention_mask=inputs["attention_mask"].to(device), max_new_tokens=600)
32output_text = tokenizer.decode(output[0])
33answer = output_text.split("assistant<|end_of_role|>")[1]
34print("Answer: " + answer)
35
36# Generate certainty score
37uq_generation_prompt = "<|start_of_role|>certainty<|end_of_role|>"
38uq_chat = [
39 {
40 "role": "system",
41 "content": ""
42 },
43 {
44 "role": "user",
45 "content": question
46 },
47 {
48 "role": "assistant",
49 "content": answer
50 },
51]
52
53uq_text = tokenizer.apply_chat_template(uq_chat,tokenize=False) + uq_generation_prompt
54# remove automatic system prompt
55string_to_remove = tokenizer.apply_chat_template(uq_chat[0:1], tokenize=False,add_generation_prompt=False)
56input_text = input_text[len(string_to_remove):]
57uq_text = uq_text[len(string_to_remove):]
58
59# tokenize and generate
60inputs = tokenizer(uq_text, return_tensors="pt")
61output = model_UQ.generate(inputs["input_ids"].to(device), attention_mask=inputs["attention_mask"].to(device), max_new_tokens=1)
62output_text = tokenizer.decode(output[0])
63uq_score = int(output_text[-1])
64print("Certainty: " + str(5 + uq_score * 10) + "%")
A.3 Hallucination Detection Quickstart Example
As explained in Section 4, it is highly recommended to use the Hallucination Detection model through Granite IO. However, if you prefer to invoke the model directly, you can use the following code.
1
2import torch
3from transformers import AutoTokenizer, AutoModelForCausalLM
4from peft import PeftModel, PeftConfig
5from nltk import tokenize
6import json
7
8BASE_NAME = "ibm-granite/granite-3.2-8b-instruct"
9LORA_NAME = "ibm-granite/granite-3.2-8b-lora-rag-hallucination-detection"
10device=torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’)
11
12tokenizer = AutoTokenizer.from_pretrained(BASE_NAME, padding_side=’left’, trust_remote_code=True)
13model_base = AutoModelForCausalLM.from_pretrained(BASE_NAME, device_map="auto")
14model_hallucination = PeftModel.from_pretrained(model_base, LORA_NAME)
15
16hallucination_sys_prompt = "Split the last assistant response into individual sentences. For each sentence in the last assistant response, identify the faithfulness score range. Ensure that your output includes all response sentence IDs, and for each response sentence ID, provide the corresponding faithfulness score range. The output must be a json structure."
17
18def format_conversation(conversation):
19 response_sents = tokenize.sent_tokenize(conversation[-1]["content"])
20 response_sents_with_ids = []
21 for ind, sent in enumerate(response_sents):
22 response_sents_with_ids.append(f"<r{ind}> {sent}")
23 conversation[-1]["content"] = .join(response_sents_with_ids)
24 conversation.append({
25 "role": "system",
26 "content": hallucination_sys_prompt
27 })
28 return conversation
29
30
31conversation = [
32 {
33 "role": "user",
34 "content": "What happened to Dennis Wilson of the Beach Boys in 1983?"
35 },
36 {
37 "role": "assistant",
38 "content": "Dennis Wilson of the Beach Boys drowned in Marina del Rey on December 28, 1983, while diving from a friend’s boat trying to recover items that he had previously thrown overboard in fits of rage. Forensic pathologists believed that Dennis experienced shallow-water blackout just before his death"
39 }
40]
41input_conversation = format_conversation(conversation=conversation)
42
43documents = [
44 {
45 "doc_id": 1,
46 "text": "The Beach Boys are an American rock band formed in Hawthorne, California, in 1961. The group’s original lineup consisted of brothers Brian, Dennis, and Carl Wilson; their cousin Mike Love; and their friend Al Jardine. Distinguished by their vocal harmonies and early surf songs, they are one of the most influential acts of the rock era. The band drew on the music of jazz-based vocal groups, 1950s rock and roll, and black R&B to create their unique sound, and with Brian as composer, arranger, producer, and de facto leader, often incorporated classical or jazz elements and unconventional recording techniques in innovative ways. In 1983, tensions between Dennis and Love escalated so high that each obtained a restraining order against each other. With the rest of the band fearing that he would end up like Brian, Dennis was given an ultimatum after his last performance in November 1983 to check into rehab for his alcohol problems or be banned from performing live with them. Dennis checked into rehab for his chance to get sober, but on December 28, 1983, he fatally drowned in Marina del Rey while diving from a friend’s boat trying to recover items that he had previously thrown overboard in fits of rage."
47 },
48 {
49 "doc_id": 2,
50 "text": "A cigarette smoker since the age of 13, Carl was diagnosed with lung cancer after becoming ill at his vacation home in Hawaii, in early 1997. Despite his illness, Carl continued to perform while undergoing chemotherapy. He played and sang throughout the Beach Boys’ entire summer tour which ended in the fall of 1997. During the performances, he sat on a stool, but he stood while singing \"God Only Knows\". Carl died of lung cancer in Los Angeles, surrounded by his family, on February 6, 1998, just two months after the death of his mother, Audree Wilson. He was interred at Westwood Village Memorial Park Cemetery in Los Angeles."
51 },
52 {
53 "doc_id": 3,
54 "text": "Carl Dean Wilson (December 21, 1946 - February 6, 1998) was an American musician, singer, and songwriter who co-founded the Beach Boys. He is best remembered as their lead guitarist, as the youngest brother of bandmates Brian and Dennis Wilson, and as the group’s de facto leader in the early 1970s. He was also the band’s musical director on stage from 1965 until his death. Influenced by the guitar playing of Chuck Berry and the Ventures, Carl’s initial role in the group was that of lead guitarist and backing vocals, but he performed lead vocals on several of their later hits, including \"God Only Knows\" (1966), \"Good Vibrations\" (1966), and \"Kokomo\" (1988). By the early 1980s the Beach Boys were in disarray; the band had split into several camps. Frustrated with the band’s sluggishness to record new material and reluctance to rehearse, Wilson took a leave of absence in 1981. He quickly recorded and released a solo album, Carl Wilson, composed largely of rock n’ roll songs co-written with Myrna Smith-Schilling, a former backing vocalist for Elvis Presley and Aretha Franklin, and wife of Wilson’s then-manager Jerry Schilling. The album briefly charted, and its second single, \"Heaven\", reached the top 20 on Billboard’s Adult Contemporary chart."
55 }
56]
57
58# Generate answer
59input_text = tokenizer.apply_chat_template(conversation=input_conversation, documents=documents, tokenize=False)
60
61inputs = tokenizer(input_text, return_tensors="pt")
62output = model_hallucination.generate(inputs["input_ids"].to(device), attention_mask=inputs["attention_mask"].to(device), max_new_tokens=500)
63output_text = tokenizer.decode(output[0][inputs["input_ids"].shape[1]:], skip_special_tokens=True)
64print("Output: " + json.loads(output_text))
A.4 Answerability Determination Quickstart Example
The following code describes how to use the Answerability Determination model.
1
2import torch
3from transformers import AutoTokenizer, AutoModelForCausalLM
4from peft import PeftModel
5from peft import PeftModelForCausalLM as lora_model
6
7device=torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’)
8
9ANSWERABILITY_PROMPT = "<|start_of_role|>answerability<|end_of_role|>"
10BASE_NAME = "ibm-granite/granite-3.2-8b-instruct"
11LORA_NAME = "ibm-granite/granite-3.2-8b-lora-rag-answerability-prediction"
12
13tokenizer = AutoTokenizer.from_pretrained(BASE_NAME, padding_side=’left’,trust_remote_code=True)
14model_base = AutoModelForCausalLM.from_pretrained(BASE_NAME,device_map="auto")
15model_answerability = PeftModel.from_pretrained(model_base, LORA_NAME)
16
17convo = [
18 {"role": "user", "content": "name the viceroy at the time of quit india movement"},
19 {"role": "assistant", "content": "The Viceroy at the time of the Quit India Movement was Lord Linlithgow."},
20 {"role": "user", "content": "opposition leader for the movement"},
21 {"role": "assistant", "content": "A movement was launched at the Bombay session of the All - India Congress Committee by Mahatma Gandhi on 8 August 1942. Many students paid attention to Subhas Chandra Bose as a leader, who was in exile. He organized the Indian Legion in Germany , reorganized the Indian National Army with Japanese assistance , and soliciting help from the Axis Powers , conducted a guerrilla war against the British authorities."},
22 {"role": "user", "content": "who is Patel?"},
23 {"role": "assistant", "content": "Vallabhbhai Patel, popularly known as Sardar Patel , was the first Deputy Prime Minister of India . He was an Indian barrister and statesman , a senior leader of the Indian National Congress and a founding father of the Republic of India who played a leading role in the country ’s struggle for independence and guided its integration into a united , independent nation . He acted as de facto Supreme Commander - in - chief of Indian army during the political integration of India and the Indo - Pakistani War of 1947 . He is credited by historians with ensuring the success of the rebellion across India. He was arrested and was imprisoned with the entire Congress Working Committee from 1942 to 1945"},
24 {"role": "user", "content": "how do you pronounce Vallabhbhai?"},
25 ]
26
27
28documents = [
[…]
32]
33
34convo = [{"role":"system", "content": ""}] +convo
35
36string = tokenizer.apply_chat_template(convo,documents=documents, tokenize=False,add_generation_prompt=False)
37string_to_remove = tokenizer.apply_chat_template(convo[0:1], tokenize=False,add_generation_prompt=False)
38string = string[len(string_to_remove):]
39inputs = string + ANSWERABILITY_PROMPT
40
41inputT = tokenizer(inputs, return_tensors="pt")
42
43output = model_answerability.generate(inputT["input_ids"].to(device), attention_mask=inputT["attention_mask"].to(device), max_new_tokens=3)
44output_text = tokenizer.decode(output[0])
45answer = output_text.split(ANSWERABILITY_PROMPT)[1]
46print(answer)
A.5 Citation Generation Quickstart Example
As explained in Section 6, it is highly recommended to use the Citation Generation model through Granite IO. However, if you prefer to invoke the model directly, you can use the following code. Note that the code assumes that the documents and the last assistant response have been already split into sentences.
1import torch
2from transformers import AutoTokenizer, AutoModelForCausalLM
3from peft import PeftModel, PeftConfig
4import json
5
6BASE_NAME = "ibm-granite/granite-3.2-8b-instruct"
7LORA_NAME = "ibm-granite/granite-3.2-8b-lora-rag-citation-generation"
8device=torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’)
9
10tokenizer = AutoTokenizer.from_pretrained(BASE_NAME, padding_side=’left’, trust_remote_code=True)
11model_base = AutoModelForCausalLM.from_pretrained(BASE_NAME, device_map="auto")
12model_citation = PeftModel.from_pretrained(model_base, LORA_NAME)
13
14conversation = [
15 {"role": "user", "content": "What is the visibility level of Git Repos and Issue Tracking projects?"},
16 {"role": "assistant", "content": "<r0> Git Repos and Issue Tracking projects can have one of the following visibility levels: private, internal, or public. <r1> Private projects are visible only to project members, internal projects are visible to all users that are logged in to IBM Cloud, and public projects are visible to anyone. <r2> By default, new projects are set to private visibility level, which is the most secure for your data."}]
17
18documents = [
[…]
21
22# Add system prompt
23citation_sys_prompt = "Split the last assistant response into individual sentences. For each sentence in the response, identify the statement IDs from the documents that it references. Ensure that your output includes all response sentence IDs, and for each response sentence ID, provide the corresponding referring document sentence IDs."
24conversation.append({"role": "system", "content": citation_sys_prompt})
25
26# Generate answer
27input_text = tokenizer.apply_chat_template(conversation=conversation, documents=documents, tokenize=False)
28inputs = tokenizer(input_text, return_tensors="pt")
29output = model_citation.generate(inputs["input_ids"].to(device), attention_mask=inputs["attention_mask"].to(device), max_new_tokens=500)
30output_text = tokenizer.decode(output[0][inputs["input_ids"].shape[1]:], skip_special_tokens=True)
31print("Output: ")
32print(json.loads(output_text))

### v2

@@ Appendix A Technical Appendices and Supplementary Material
A.1 Changes in performance from Granite Instruct 3.2 to 3.3
In this section we compare performance for intrinsics fine-tuned on Granite 3.3 8b Instruct to earlier versions trained on Granite 3.2 8b Instruct (we found no significant difference in performance for the Query Rewrite, Hallucination Detection, and Uncertainty Quantification intrinsics.)
A.1.1 Answerability Determination
We compare the performance of the answerability determination intrinsic fine-tuned on Granite 3.3 8b Instruct to an earlier version trained on Granite 3.2 8b Instruct. Table 23 and Table 24 presents a detailed evaluation on the MT-RAG and SQUAD datasets respectively, reporting precision, recall, and F1 scores for both answerable and unanswerable queries. Compared to Granite 3.2, the Granite 3.3-based model shows improved recall on answerable examples, but this comes at the expense of performance on unanswerable queries. Aggregate performance remains quite comparable across models, though Granite 3.3 performs slightly worse on SQUAD.
unanswerable answerable W.F1
P R F1 P R F1
Granite 3.2 8b
Instruct LoRA
89.1 93.7 91.3 92.3 86.7 89.4 90.5
Granite 3.3 8b
Instruct LoRA 89.8 91.8 90.8 90.3 87.9 89.1 90
Table 23: Evaluation of the Granite 3.2 and 3.3 versions of the answerability classification intrinsic on MT-RAG.
unanswerable answerable W.F1
P R F1 P R F1
Granite 3.2 8b
Instruct LoRA
84.2 68 75.2 73.1 87.2 79.5 77.4
Granite 3.3 8b
Instruct LoRA 88.1 59.3 70.9 69.3 92 79 75
Table 24: Evaluation of the Granite 3.2 and 3.3 versions of the answerability classification intrinsic on SQUAD.
A.1.2 Citation Generation
We compare the performance of the latest citation generation intrinsic trained on IBM Granite 3.3 8b instruct to an earlier version of the intrinsic trained on the IBM Granite 3.2 8b instruct model. Tables 25 and 26 show the performance of the two versions of the intrinsic compared to the baselines when evaluated on ALCE and LongBench-Cite, respectively. The experimental settings are identical to the ones described in Sections 9.2.1 and 9.2.2 for ALCE and LongBench-Cite, respectively.
Llama-3.1-70B-Instruct Llama-3.1-70B-Instruct 61.4 58.1 59.7
Llama-3.1-70B-Instruct Granite-3.3-8B LoRA citations 55.4 65.0 59.8
Mixtral-8x22B-Instruct Mixtral-8x22B-Instruct 62.2 62.5 62.3
Mixtral-8x22B-Instruct Granite-3.3-8B LoRA citations 55.6 69.0 61.6
Table 25: Evaluation of the Granite 3.2 and 3.3 versions of the citation generation intrinsic on ALCE
Model generating response Model generating citations Longbench-Chat (en) MultifieldQA (en) HotpotQA GovReport AVG
R P F1 R P F1 R P F1 R P F1 F1
Llama-3.1-70B-Instruct Llama-3.1-70B-Instruct 27.0 34.4 26.1 46.1 63.3 49.7 34.0 39.4 30.2 55.0 77.5 62.0 42.0
Llama-3.1-70B-Instruct Granite-3.2-8B LoRA citations 61.9 68.6 62.0 71.2 84.1 74.3 66.8 73.3 65.4 70.3 83.6 75.4 69.3
Llama-3.1-70B-Instruct Granite-3.3-8B LoRA citations 57.6 60.3 58.4 71.5 82.5 75.0 65.3 71.3 63.8 72.8 83.5 77.2 68.6
Table 26: Evaluation of the Granite 3.2 and 3.3 versions of the citation generation intrinsic on LongBench-Cite
We observe that the two versions of the citations generation intrinsic exhibit similar performance. When evaluated on ALCE, the intrinsic trained on IBM Granite 3.3 8b instruct performs the same or slightly better than the one trained on 3.2, while when evaluated on LongBench-Cite, the 3.3 version is overall slightly worse than the 3.2 version. However, the difference is pretty small with the overall performance of the two versions being in all conducted experiments within 0.7 percentage points of each other.